# Towards Robust Stability Prediction in Smart Grids: GAN-based Approach under Data Constraints and Adversarial Challenges

### v1

@@ IV Reference Attacks
In the white-box scenario, an attacker can exploit various state-of-the-art adversarial attacks. While numerous attacks exist, most have been evaluated primarily in multi-classification tasks and are not specifically designed for binary classification tasks like ours. We concentrate on specific attacks recognized for their effectiveness in exposing weaknesses, particularly in the context of decision-making in smart grids, as supported by existing literature [17, 3].
@@
In the grey-box scenario, the same adversarial attacks are employed; however, adversaries only have access to genuine data and a surrogate model. They utilize this surrogate model to generate adversarial data, which is then tested against the primary model. This method assesses the robustness of the primary model without granting direct access to it.
@@
The equation maximizes the loss $L$ between the model’s prediction $f(x+\epsilon)$ and the true label $y$ while constraining the perturbation $\epsilon$ within a specified norm limit $\|\epsilon\|_{p}\leq\gamma$.
@@
V GAN-Stability: Our Proposed Stability Prediction System
In this section, we present our proposed stability prediction system, whose architecture is summarized in Figure 1. In particular, training of the discriminator is done employing stable data ( 1), synthetic unstable data created by the generator starting from noise ( 2), and adversarial samples generated by applying attacks to stable data ( 3).
@@ V-A Architecture
Our architecture leverages a GAN model for smart grid stability prediction, where the generator creates synthetic data representing potential instances of instability, addressing the challenge of acquiring rare unstable data in real-world scenarios. The discriminator learns to differentiate between real (stable) data and synthetic (unstable) data, enhancing its ability to detect anomalies and predict stability accurately even when we do not have access to one label (unstable) in the training data. In this context, the generator does not have direct access to real data; its learning process relies solely on its interaction with the discriminator, which has access to both generated and real samples.
The generator model, deliberately simpler than the discriminator, consists of four fully connected layers, with the number of neurons ranging from 12 (representing the number of features) to 128. The discriminator, instead, comprises five fully connected layers, with the neuron count ranging from 12 to 512. The specific details of our GAN-based stability prediction system architecture can be seen in Table I.
TABLE I: our GAN-based stability prediction system Architecture
@@ V-B Training
In our approach, we utilize a specialized training procedure for GAN models. Here, the generator begins with random noise and attempts to deceive the discriminator by generating synthetic data samples that resemble genuine instances. Concurrently, the discriminator is trained on real data, which, in our case, comprises just stable instances of the dataset, as well as on the fake data generated by the generator. Over time, the discriminator learns to distinguish between genuine stable data and the synthetic data produced by the generator, which may represent potential instances of instability lying outside the distribution of stable labels.
Unlike the traditional usage of GANs, where the goal is for the generator to eventually converge and produce data that closely resembles the real data such that $p_{g}(x)\approx p_{\text{data}}(x)$, we intentionally prevent the generator from reaching that stage of convergence. The optimization process for the generator is intentionally constrained to prevent it from reaching convergence. This modified min-max problem reflects our novel approach where the generator’s role is not to reach the standard equilibrium, but to guide the discriminator into learning better classification boundaries for stable and unstable instances.
@@
$\text{ReLU}(m-\|x-s\|)=\begin{cases}m-\|x-s\|,&\text{if }\|x-s\|<m,\\ 0,&\text{if }\|x-s\|\geq m.\end{cases}$
@@
$L_{G}=-\mathbb{E}_{z\sim p_{z}}\left[\log D(G(z))\right]+\\ \mathbb{E}_{x\sim p_{g},\,s\sim p_{\text{stable}}}\left[\text{ReLU}(m-\|x-s\|)\right],$
@@
The choice of a margin $m=4$ is justified based on the distribution of features in the dataset and their deviations. Specifically, the margin is selected to ensure that the generated samples $p_{g}$ maintain a safe distance from the stable class $p_{\text{stable}}$, avoiding overlap while still exploring plausible regions of the data space. By analyzing the scale and standard deviations of the feature values, a margin of 4 represents a sufficient threshold to distinguish generated samples from the stable class while not overly restricting the generator’s ability to explore diverse, non-stable regions. The discriminator’s ability to classify stability and instability does not rely on the generator producing perfectly unstable data. Instead, the generator serves as an adversarial tool that pushes the discriminator to refine its decision boundaries. By doing so, the discriminator learns to focus on feature combinations that determine stability or instability, regardless of whether the generator’s outputs match real-world patterns.
@@
As depicted in Figure 2, the generator’s loss begins at a high value and initially decreases as the generator starts challenging the discriminator. This early improvement highlights the generator’s growing ability to deceive a discriminator that is still in the very early stages of learning, having been trained for less than 50 epochs. During this phase, the repulsion loss (shown in Figure 3) starts with high values, reflecting the generator’s proximity to the real data distribution. However, the repulsion loss subsequently becomes optimized, enforcing divergence from the stable data distribution. This process, coupled with the discriminator’s growing advantage, results in a gradual increase in the generator’s loss, eventually reaching values between 10 and 40. The increasing generator loss reflects the escalating difficulty of producing samples that can deceive the discriminator.
Initially, the repulsion loss has higher values. However, as training progresses and the discriminator gains the upper hand, the repulsion loss is optimized to push the generator’s output away from stable samples. This effect is evidenced by the decreasing repulsion loss values during later epochs, corresponding to the generator’s divergence from stability.
The discriminator’s architectural advantage, with its deeper network design, further reinforces this dynamic. Its capacity to learn complex representations allows it to maintain dominance throughout training, as seen by its relatively stable loss values in Figure 2. This structural superiority, coupled with prolonged training and repulsion loss, ensures the generator does not converge too closely to the stable data distribution. Instead, it generates data that is more representative of potential instabilities, providing a more effective adversarial challenge and enhancing the model’s robustness.
Through the interplay of the repulsion term, extended training, and the discriminator’s architecture, our approach fosters a dynamic adversarial process. This controlled imbalance ensures the generator continually adapts, producing harder-to-fool samples that represent potential instabilities more effectively.
Despite the absence of direct unstable examples in the training set, the discriminator effectively generalizes its understanding of stability and instability based on the provided data. This enhances the model’s robustness in predicting system stability. GAN-Stability is trained over 250 epochs with a learning rate of 0.0002 and a batch size of 4, ensuring thorough learning and refinement of the classification boundaries.
V-C Adversarial Training
Together with real stable data ( 1 in Figure 1) and the samples crafted by the generator ( 2), we introduce another source of data ( 3) composed by a novel adversarial layer which is a novelty respect the traditional GAN training process. The detailed steps about our proposed training approach can be found in Algorithm 1. Initially, the training steps follow the standard procedure as mentioned before, commencing with training the discriminator on real data labeled as stable ( 1). Upon backpropagating the discriminator’s loss, we proceed to generate fake data using the generator ( 2), which serves as our potential instances of instability in the grid. This generation process begins by initializing with a random tensor of latent inputs, which the generator model then processes to generate samples. Subsequently, we evaluate these generated samples using the discriminator and backpropagate the loss accordingly.
@@ VI-A Dataset
VI-B Baseline Evaluation
In the evaluation stage, we begin by establishing the baseline performance of our GAN-based stability prediction systems. This assessment is conducted before integrating any countermeasures, such as our novel adversarial training layer, and before exposing the system to adversarial attacks. Initially, we train our proposed model using the available training data (90% of stable class with no samples from unstable class as discussed in Section VI-A). Following the training phase, we proceed to evaluate the effectiveness of our GAN-based stability prediction system on the test dataset which compromises 10% of stable instances and all samples from the unstable class.
@@ VI-B Baseline Evaluation
The discriminator’s effective generalization without access to the unstable class in training and despite the generator’s increasing loss indicates that the extended training process, the discriminator’s deeper architecture and the repulsion loss exposed it to different data from the stable class, enabling it to accurately classify unseen unstable instances. By allowing the discriminator to outperform the generator during training, adding the repulsion loss, and incorporating more layers and neurons, the adversarial process enhanced its ability to identify deviations from the stable class. This strong performance suggests that even without access to unstable class in the training, the diverse samples that the generator produces are sufficient for the discriminator to learn the complex characteristics of instability.
VI-C GAN-stability with Adversarial Training Evaluation
In this section, we evaluate our GAN-based stability prediction system, augmented with our novel adversarial training layer as discussed in Section V-C. To assess the robustness of our model against adversarial attacks discussed in Section IV, we utilize the TorchAttacks and Adversarial Robustness Toolbox (ART) library for attack implementation. We subject the discriminator model and testing dataset to various attacks, including FGSM, RFGSM, BIM, and PGD, as detailed in Section IV. We employ an epsilon value of 0.05 for each attack, which signifies the strength of the attack and the magnitude of perturbation added to the data. The selection of 0.05 strikes a balance between the attack’s power and the model’s susceptibility, effectively challenging our model’s ability to detect perturbations introduced by the attack. This value allows us to explore the attack’s effectiveness while ensuring it remains within manageable bounds, thus facilitating a comprehensive evaluation of the model’s robustness. By constraining the attack within manageable bounds, we mitigate potential risks such as susceptibility to ADS or expert human intervention in a real world scenario of a possible attack. In addition, we train the surrogate model, as previously described in Section III-B, utilizing 70% of the entire dataset (both labels) for training.
@@ VI-C GAN-stability with Adversarial Training Evaluation
Our evaluation shows the effectiveness of our stability prediction system in accurately classifying state-of-the-art whitebox adversarial attacks. With a mean accuracy of 0.989, our system adeptly identifies these attacks as belonging to the unstable class. Furthermore, in the greybox setting, where attacks are launched against our stability prediction system using adversarial samples generated from a surrogate model, our system achieves a mean accuracy of 0.954. Detailed findings are presented in Table II. By accurately identifying these attacks as instances of instability, our stability prediction model can promptly raise an alarm to grid operators, enabling them to investigate the situation and take necessary precautions to prevent potential problems caused by the attack. This proactive approach reduces the necessity for employing another model acting as an ADS, which could introduce latency or overload to the grid communication system.
TABLE II: Accuracy of GAN-Stability in classifying attacks as unstable samples.
@@
FGSM BIM RFGSM PGD
Whitebox $1.0$ $0.957$ $1.0$ $1.0$
Greybox $0.954$ $0.955$ $0.954$ $0.956$
Following the augmentation of GAN-Stability with the new adversarial training layer, it becomes essential to re-evaluate its performance against the main task of stability prediction to ensure that the augmentation does not compromise the efficacy of the system. In these settings, our system achieves a mean accuracy of 0.975 considering both classes. Specifically, when tested against stable instances of the test set, our model achieves an accuracy of 0.912, and when tested against unstable instances, it achieves an accuracy of 0.979. The integration of the adversarial attack layer significantly boosts the performance of the discriminator, despite the small deviations introduced by the attack. By incorporating adversarial samples, the discriminator is exposed to subtle perturbations of real stable data. These adversarial examples act as challenging cases, forcing the discriminator to become more sensitive to nuanced changes in the input data. Although these perturbations may seem minor, they encourage the discriminator to refine its decision boundary and strengthen its ability to differentiate between truly stable and unstable instances. The combined impact of the repulsion loss, the discriminator’s architectural superiority, the extended training period, and the adversarial training layer ensures the system is equipped to accurately identify instability while maintaining strong generalization across diverse data distributions. The summary of the results is presented in Table III.
TABLE III: Accuracy and F1 score of GAN-Stability with and without Adversarial Training (AT).
@@
0.975 0.987
VI-D Benchmarking Against State-of-the-Art
In this section, we compare our proposed system with state-of-the-art models in the literature, utilizing the same dataset. As indicated in Table IV, many of the existing models achieve high accuracies. However, it is important to consider that these models have been trained on the entire dataset with access to both labels, which may not be feasible in real-world scenarios. Collecting instances of instability requires significant time, resources, and human expertise. Importantly, it is almost not possible to train these models with only one label, highlighting a key advantage of our proposed approach. Additionally, it is noteworthy that none of the models in the literature address the potential threat of adversarial attacks against these systems or implement countermeasures to mitigate such attacks. Therefore, it is reasonable to assume that many of these systems would be vulnerable to adversarial attacks. Unlike these models, our proposed approach is trained with only one label, making it more practical for real-world deployment. Furthermore, by incorporating adversarial training layers, we have enhanced the resilience of our system against adversarial attacks, a consideration often overlooked in previous works.
TABLE IV: Comparison of GAN-Stability with other state-of-the-art models. Only our system considers adversarial attacks.

### v2

@@ 4 Reference Attacks
In the white-box scenario, an attacker can exploit various state-of-the-art adversarial attacks. While many such attacks have been proposed, most have been evaluated primarily in multi-class classification tasks and are not specifically tailored for binary classification problems like ours. We focus on a subset of attacks known for their effectiveness in revealing model vulnerabilities, particularly in decision-making contexts within smart grids, as supported by existing literature [17, 3].
@@
In the grey-box scenarios, the same adversarial attacks are employed. However, in the first grey-box scenario, adversaries only have access to genuine data and a surrogate model. They utilize this surrogate model to generate adversarial data, which is then tested against the primary model. This method assesses the robustness of the primary model without granting direct access to it.
We will also consider another attack scenario (grey-box 2) called GAN-GRID [3], a more sophisticated generative attack targeting the stability prediction system. This attack does not require access to the data or model architecture; it only needs the ability to query the model to craft adversarial inputs that can be classified as stable instances. These deceptive inputs could then be injected into the grid to mislead the stability prediction system.
@@
Equation 1 maximizes the loss $L$ between the model’s prediction $f(x+\epsilon)$ and the true label $y$ while constraining the perturbation $\epsilon$ within a specified norm limit $\|\epsilon\|_{p}\leq\gamma$.
@@
  • 5.
    GAN-GRID: is an adversarial attack that leverages the generator network of a GAN to create adversarial samples. The generator is trained using reinforcement learning methods, with a grid stability prediction system acting as a fixed discriminator or oracle. The training process focuses on refining the generator’s ability to produce effective adversarial samples based on feedback from the stability prediction system. In each training episode, the generator generates a sample from the latent space, which is then assessed by the stability prediction model. The model assigns a reward based on how closely the generated sample aligns with the target predictions, guiding the generator’s optimization process [3].
5 GAN-Stability: Our Proposed Stability Prediction System
In this section, we present our proposed stability prediction system, whose architecture is summarized in Figure 1. In particular, training of the discriminator is done employing stable data ( 1), synthetic Out-Of-Distribution (OOD) data relative to the stable class, generated by the generator from noise ( 2), and adversarial samples generated by applying attacks to stable data ( 3).
@@ 5.1 Architecture
Our architecture employs a GAN-based framework for smart grid stability prediction, where the generator produces synthetic OOD samples relative to the stable class. These OOD samples may include measurement anomalies, fault-induced disturbances, instability events, or other deviations that do not conform to the stable data distribution.
The discriminator is trained to distinguish between real (stable) data and these synthetic OOD samples. By learning to differentiate stable conditions from any form of unstable behavior—without explicitly requiring labeled unstable data—our framework enhances anomaly detection and improves stability prediction. This approach enables the model to develop a more robust decision boundary, allowing for more accurate and generalized assessments of grid stability.
In this context, the generator does not have direct access to real data; its learning process relies solely on its interaction with the discriminator, which has access to both generated and real samples. The generator model, deliberately simpler than the discriminator, consists of four fully connected layers, with the number of neurons ranging from 12 (representing the number of features) to 128. The discriminator, instead, comprises five fully connected layers, with the neuron count ranging from 12 to 512. The specific details of our GAN-based stability prediction system architecture can be seen in Table 1.
Table 1: The architecture of our GAN-based stability prediction system.
@@ 5.2 Training
In our approach, we employ a specialized training procedure for GAN models. The generator starts with random noise and aims to challenge the discriminator by producing synthetic data samples. However, rather than generating near-realistic data, the generator is specifically designed to create OOD samples relative to the stable class.
Meanwhile, the discriminator is trained on real data—comprising only stable instances from the dataset—alongside the synthetic OOD samples generated by the generator. Over time, the discriminator learns to differentiate between genuine stable data and these synthetic samples, effectively identifying deviations that fall outside the distribution of stable conditions.
Unlike the traditional usage of GANs, where the goal is for the generator to eventually converge and produce data that closely resembles the real data such that $p_{g}(x)\approx p_{\text{data}}(x)$, we intentionally prevent the generator from reaching that stage of convergence. The optimization process for the generator is deliberately constrained to prevent it from fully converging. This modified min-max formulation aligns with our novel approach, where the generator’s objective is not to reach the typical GAN equilibrium but rather to help the discriminator learn more effective classification boundaries between stable data and OOD samples which could potentially involve unstable instances.
@@
$\text{ReLU}(m-\|x-s\|)=\begin{cases}m-\|x-s\|,&\text{if }\|x-s\|<m,\\ 0,&\text{if }\|x-s\|\geq m.\end{cases}$ (6)
@@
$L_{G}=-\mathbb{E}_{z\sim p_{z}}\left[\log D(G(z))\right]+\mathbb{E}_{x\sim p_{g},\,s\sim p_{\text{stable}}}\left[\text{ReLU}(m-\|x-s\|)\right],$ (7)
@@
Our objective is not for the generator to converge to the stable class but rather to guide it toward generating data that moves closer to the stable class distribution while maintaining a safe margin enforced by the repulsion loss. Since the latent distribution $p_{z}$ (e.g., a Gaussian) has full support over $\mathbb{R}^{d}$, the generator inherently possesses the capacity to map to any region in the data space $X$. This means it is not restricted to reconstruct only the stable class but can generate samples in diverse and potentially unstable regions. To guide the generator away from the stable distribution $p_{\text{stable}}$, we incorporate a repulsion loss that acts as a soft constraint. This loss creates an exclusion zone around the stable set $S\subset X$, encouraging the generator to produce samples at least a margin $m$ away from it. Formally, the repulsion loss drives generated samples into the set $X_{\text{OOD}}\subset X\setminus B_{m}(S)$, defined as:
$X_{\text{OOD}}:=\left\{x\in X:\|x-s\|\geq m\text{ for all }s\in S\right\},$
where $B_{m}(S)$ denotes the $m$-ball around the stable set. As training progresses, the distribution $p_{g}$ of generated samples converges to high-density regions within this complement, effectively modeling plausible but non-stable areas in the data space. At the same time, adversarial pressure from the discriminator ensures these samples remain realistic, as it continues to evaluate whether $G(z)$ resembles true data. This dynamic maintains the generator’s outputs on the data manifold while steering them outside the stable cluster, supporting the overall objective of improving discriminator boundaries between stable and OOD regions.
The choice of the margin m is grounded in the statistical characteristics of the stable class and the normalization applied to the dataset. After applying z-score normalization, each feature in the dataset has a standard deviation of approximately 1. Analysis of the stable class distribution showed that the majority of samples lie within three standard deviations from the mean , this follows the Empirical Rule, which states that approximately 99.7% of values in a normal distribution fall within three standard deviations from the mean . To provide a conservative buffer and encourage exploration beyond this high-density region, we selected a margin of four standard deviations. This corresponds to $m=4$ in the normalized space. This value was empirically validated to ensure that the generated samples $p_{g}$ maintain a safe distance from the stable class $p_{\text{stable}}$, avoiding overlap while still exploring plausible regions of the data space.
The discriminator’s ability to classify stable and unstable samples does not rely on the generator producing perfectly unstable real-world data. Instead, the generator acts as an adversarial tool, challenging the discriminator to refine its decision boundaries. Through this process, the discriminator learns to focus on key feature combinations that define stability, rather than relying on direct comparisons to real-world patterns. The generated OOD samples may encompass measurement anomalies, fault-induced disturbances, or other deviations that fall outside the stable data distribution. This approach enables the discriminator to generalize better, improving its ability to detect atypical scenarios without requiring explicit labels for unstable conditions.
@@
As shown in Figure 2, the generator’s loss starts at a high value and initially decreases as the generator begins to effectively challenge the discriminator. This early improvement highlights the generator’s growing ability to deceive a discriminator that is still in the very early stages of learning, having been trained for less than 50 epochs. During this phase, the repulsion loss (Figure 3) starts with high values, reflecting the generator’s proximity to the real data distribution. As training progresses, the repulsion loss decreases, indicating successful optimization that drives the generator to diverge from the stable data distribution. This process, coupled with the discriminator’s growing advantage, results in a gradual increase in the generator’s loss, eventually reaching values between 10 and 40.
The discriminator’s architectural advantage, with its deeper network design, further reinforces this dynamic. Its capacity to learn complex representations allows it to maintain dominance throughout training, as seen by its relatively stable loss values in Figure 2. This structural superiority, coupled with prolonged training and repulsion loss, ensures the generator does not converge too closely to the stable data distribution. Instead, it generates OOD samples with respect to the stable class, which may be representative of any form of unstable behavior.
GAN-Stability is trained over 250 epochs with a learning rate of 0.0002 and a batch size of 4, ensuring thorough learning and refinement of the classification boundaries.
5.3 Adversarial Training
Together with real stable data ( 1 in Figure 1) and the samples crafted by the generator ( 2), we introduce another source of data ( 3) composed by a novel adversarial layer which is a novelty respect the traditional GAN training process. The detailed steps about our proposed training approach can be found in Algorithm 1. Initially, the training steps follow the standard procedure as mentioned before, commencing with training the discriminator on real data labeled as stable ( 1). Upon backpropagating the discriminator’s loss, we proceed to generate fake data using the generator ( 2), This generation process begins by initializing with a random tensor of latent inputs, which the generator model then processes to generate samples. Subsequently, we evaluate these generated samples using the discriminator and backpropagate the loss accordingly.
@@ 6.1 Dataset
6.2 Baseline Evaluation
In the evaluation stage, we begin by establishing the baseline performance of our GAN-based stability prediction systems. This assessment is conducted before integrating any countermeasures, such as our novel adversarial training layer, and before exposing the system to adversarial attacks. Initially, we train our proposed model using the available training data (90% of stable class with no samples from unstable class as discussed in Section 6.1). Following the training phase, we proceed to evaluate the effectiveness of our GAN-based stability prediction system on the test dataset which compromises 10% of stable instances and all samples from the unstable class.
@@ 6.2 Baseline Evaluation
By allowing the discriminator to outperform the generator during training, adding the repulsion loss, and incorporating more layers and neurons, the adversarial process enhanced its ability to identify deviations from the stable class. This strong performance suggests that even without access to unstable class in the training, the diverse samples that the generator produces are sufficient for the discriminator to learn the complex characteristics of unstable behavior in any form.
6.3 GAN-stability with Adversarial Training Evaluation
In this section, we evaluate our GAN-based stability prediction system, augmented with our novel adversarial training layer as discussed in Section 5.3. To assess the robustness of our model against adversarial attacks discussed in Section 4, we utilize the TorchAttacks [60] and Adversarial Robustness Toolbox (ART) [61] library for attack implementation. We subject the discriminator model and testing dataset to various attacks, including FGSM, RFGSM, BIM, PGD and GAN-GRID, as detailed in Section 4. We employ an epsilon value of 0.05 for each attack, which signifies the strength of the attack and the magnitude of perturbation added to the data. The selection of 0.05 strikes a balance between the attack’s power and the model’s susceptibility, effectively challenging our model’s ability to detect perturbations introduced by the attack. This value allows us to explore the attack’s effectiveness while ensuring it remains within manageable bounds, thus facilitating a comprehensive evaluation of the model’s robustness. By constraining the attack within manageable bounds, we mitigate potential risks such as susceptibility to ADS or expert human intervention in a real world scenario of a possible attack. In addition, we train the surrogate model, as previously described in Section 3.2, utilizing 70% of the entire dataset (both labels) for training.
@@ 6.3 GAN-stability with Adversarial Training Evaluation
Our evaluation shows the effectiveness of our stability prediction system in accurately classifying state-of-the-art white-box adversarial attacks. With a mean accuracy of 0.989, our system adeptly identifies these attacks as belonging to the unstable class. Furthermore, in the grey-box setting, where attacks are launched against our stability prediction system using adversarial samples generated from a surrogate model, our system achieves a mean accuracy of 0.988. Furthermore, in the second grey-box scenario of the GAN-GRID attack, our framework achieves an accuracy of 0.991. Detailed findings are presented in Table 2. To isolate the individual contributions of the adversarial training layer and the generator’s OOD sample generation in enhancing the discriminator’s robustness, we compare our full model against a baseline version without adversarial training. This baseline model, despite not being explicitly trained on adversarial examples, achieved a detection accuracy of 0.928 when evaluated under white-box adversarial attack conditions, 0.881 in the grey-box 1 scenario, and 0.958 in grey-box 2. These results demonstrate that the generator’s ability to produce realistic OOD samples significantly improves the model’s resilience, enabling it to distinguish between stable and unstable behaviors even without direct exposure to adversarial attacks. However, incorporating the adversarial training layer further enhances performance across all scenarios. By accurately identifying these attacks as non-stable instances, our stability prediction model can promptly raise an alarm to grid operators, enabling them to investigate the situation and take necessary precautions to prevent potential problems caused by the attack. This proactive approach reduces the need to employ another model acting as an ADS, which could introduce latency or overload to the grid communication system.
Table 2: Accuracy of GAN-Stability in classifying attacks as unstable samples. “N/A” indicates that the specific attack was not performed in that scenario.
@@
FGSM BIM RFGSM PGD GAN-GRID
White-box $1.000$ $0.957$ $1.000$ $1.000$ N/A
Grey-box 1 $0.987$ $0.988$ $0.987$ $0.989$ N/A
Grey-box 2 N/A N/A N/A N/A $0.991$
Following the augmentation of GAN-Stability with the new adversarial training layer, it becomes essential to re-evaluate its performance against the main task of stability prediction to ensure that the augmentation does not compromise the efficacy of the system. In these settings, our system achieves a mean accuracy of 0.981 considering both classes. Specifically, when tested against stable instances of the test set, our model achieves an accuracy of 0.911, and when tested against unstable instances, it achieves an accuracy of 0.986. The integration of the adversarial attack layer significantly boosts the performance of the discriminator, despite the small deviations introduced by the attack. By incorporating adversarial samples, the discriminator is exposed to subtle perturbations of real stable data. These adversarial examples act as challenging cases, forcing the discriminator to become more sensitive to nuanced changes in the input data. Although these perturbations may seem minor, they encourage the discriminator to refine its decision boundary and strengthen its ability to differentiate between truly stable and unstable instances. The combined impact of the repulsion loss, the discriminator’s architectural superiority, the extended training period, and the adversarial training layer ensures the system is equipped to accurately identify instability while maintaining strong generalization across diverse data distributions. The summary of the results is presented in Table 3. Also, as shown in Figure 4, the ROC curve illustrates the GAN-Stability’s performance across different threshold values.
Table 3: Accuracy and F1 score of GAN-Stability with and without Adversarial Training (AT).
@@
0.981 0.99
Figure 4: ROC curve for classification performance of GAN-Stability augmented with the adversarial training layer
6.4 Benchmarking Against State-of-the-Art
In this section, we compare our proposed system with state-of-the-art models from the literature using the same dataset. As shown in Table 4, many existing models achieve high accuracy. However, these models are trained on the entire dataset with access to both labels, an approach that may not be feasible in real-world scenarios due to the significant time, resources, and expertise required to collect instability instances. Notably, training such models with only one label is nearly impossible, underscoring a key advantage of our method.
Despite being trained without unstable data, our GAN-based approach exhibits superior robustness compared to some supervised models. The stability of the GAN framework facilitates effective learning without explicit exposure to unstable samples, while the generator produces OOD samples that serve as adversarial challenges, improving the discriminator’s generalization beyond the training distribution. Although certain supervised models achieve higher accuracy, they rely on labeled unstable data, whereas our model is tested on a significantly larger and more diverse dataset that includes all unstable cases. Furthermore, existing models do not address the risk of adversarial attacks or incorporate countermeasures to mitigate them. In contrast, our approach not only operates with a single-label training paradigm, making it more practical for real-world deployment, but also integrates adversarial training layers, enhancing resilience against such threats—an aspect often overlooked in prior research.
Table 4: Comparison of GAN-Stability with other state-of-the-art models. Only our system evaluates adversarial robustness, which is not reported in prior works. “N/A” indicates that adversarial robustness was not assessed in those studies.